# Conformal Survival Bands for Risk Screening under Right-Censoring

### v2

@@ 1.2 Preview of Our Contributions
For example, in Figure 1, we illustrate a scenario in which the practitioner wishes to identify “low-risk” patients—defined as those with more than an $p=80\%$ chance of surviving beyond time $t=3$ years. A patient is flagged if their conformal band lies entirely above the 80% line at $t=3$. Our method asymptotically guarantees that, among all flagged individuals, on average at least 80% will survive beyond this time point. This guarantee holds pointwise over all fixed choices of the survival probability threshold $p$ and time horizon $t$, and for “low-risk” as well as “high-risk” screening, making the method flexible and broadly applicable.
@@ 2.2 Preliminaries: Review of Conformal Inference without Censoring
If complete (uncensored) data are available, this hypotheses can be easily tested as follows. Let $\hat{S}_{T}(t\mid x)$ denote the survival function estimated by the fitted model ^$\hat{\mathcal{M}}_{T}$. We define the left-tail nonconformity scoring function as
$\hat{s}_{\mathrm{lt}}(t;x):=1-\hat{S}_{T}(t\mid x),$ (2)
which represents the predicted probability that an individual with covariates $x$ experiences an event before time $t$. This is a natural and interpretable choice of nonconformity score in this context, and—crucially—it is monotonically increasing in $t$. While our later methodology will not depend specifically on this choice of score, it does require that the scoring function $\hat{s}_{\mathrm{lt}}(t;x)$ be monotone increasing in $t$.
Using this function, we compute a non-conformity score $\hat{s}_{\mathrm{lt}}(t;X_{i})$ for each calibration point $i\in[n]$ and for the test point $X_{n+j}$. Then, following the approach of Jin and Candès (2023b), we define the left-tail conformal p-value as:
$\tilde{\phi}_{\mathrm{lt}}(t;X_{n+j}):=\frac{1+\sum_{i=1}^{n}\mathbb{I}\left\{\hat{s}_{\mathrm{lt}}(T_{i};X_{i})\geq\hat{s}_{\mathrm{lt}}(t;X_{n+j})\right\}}{n+1},$ (3)
where we assume for simplicity that all scores are almost surely distinct; this is not a limitation in practice since ties can be broken by adding some independent random noise.
@@
Assume the calibration data and test point $(X_{n+j},T_{n+j})$ are exchangeable, and the scoring function $\hat{s}_{\mathrm{lt}}(t;x)$ is monotone increasing in $t$ for all $x$. Then for any fixed $t>0$ and $\alpha\in(0,1)$, the conformal p-value $\tilde{\phi}_{\mathrm{lt}}(t;X_{n+j})$ defined in (3), computed using the full (uncensored) data, satisfies:
@@
which is useful for identifying low-risk individuals. In this case, we consider a right-tail scoring function assumed to be monotone decreasing in $t$; e.g.,
$\hat{s}_{\mathrm{rt}}(t;x):=\hat{S}_{T}(t\mid x),$ (5)
interpreted as the predicted probability of experiencing the event after time $t$. The corresponding right-tail conformal p-value is
$\tilde{\phi}_{\mathrm{rt}}(t;X_{n+j}):=\frac{1+\sum_{i=1}^{n}\mathbb{I}\left\{\hat{s}_{\mathrm{rt}}(t;X_{i})\geq\hat{s}_{\mathrm{rt}}(t;X_{n+j})\right\}}{n+1},$
@@ 2.3 Conformal Inference for Survival Analysis under Right-Censoring
$\hat{\phi}_{\mathrm{lt}}(t;X_{n+j}):=\frac{1+\sum_{i=1}^{n}E_{i}\cdot\hat{w}(T_{i},X_{i})\cdot\mathbb{I}\left\{\hat{s}_{\mathrm{lt}}(T_{i};X_{i})\geq\hat{s}_{\mathrm{lt}}(t;X_{n+j})\right\}}{1+\sum_{i=1}^{n}E_{i}\cdot\hat{w}(T_{i},X_{i})},$ (7)
where the numerator includes only data points whose true event times are observed.
@@
$\hat{\phi}_{\mathrm{rt}}(t;X_{n+j}):=\frac{1+\sum_{i=1}^{n}E_{i}\cdot\hat{w}(T_{i},X_{i})\cdot\mathbb{I}\left\{\hat{s}_{\mathrm{rt}}(T_{i};X_{i})\geq\hat{s}_{\mathrm{rt}}(t;X_{n+j})\right\}}{1+\sum_{i=1}^{n}E_{i}\cdot\hat{w}(T_{i},X_{i})}.$ (8)
@@
    • a training set of cardinality $N$ used to estimate $\hat{w}(t;x)$, $\hat{s}_{\mathrm{lt}}(t;x)$, and $\hat{s}_{\mathrm{rt}}(t;x)$;
@@
In the next section, we explain how to use IPCW conformal p-values to construct conformal survival bands.
@@ 2.4 Conformal Survival Bands for Screening High- or Low-Risk Individuals
4:  for each calibration point $(X_{i},T_{i})\in\mathcal{D}_{\text{cal}}^{\prime}$ do
5:     Compute $\hat{s}_{\mathrm{lt}}(T_{i};X_{i})$ and $\hat{s}_{\mathrm{rt}}(T_{i};X_{i})$ using Eqs. (2) and (5).
6:  end for
7:  for each time $t\in\mathcal{T}$ do
@@
9:        Compute scores $\hat{s}_{\mathrm{lt}}(t;X_{n+j})$ and $\hat{s}_{\mathrm{rt}}(t;X_{n+j})$.
@@ 2.5 Theoretical Guarantees for Calibrated Survival Bands
    For all $t>0$, the score functions $\hat{s}_{\mathrm{lt}}(t;X)$ admit a Lebesgue density $f_{t}$ satisfying
    $0<f_{\min}\leq f_{t}(u)\leq f_{\max}<\infty\quad\text{for all }u\in\mathbb{R},$
@@
Under Assumptions (A1)(A5), fix a target level $\alpha\in(0,1)$ and survival cutoff $t>0$. Let $p_{j}:=\hat{\phi}_{\mathrm{lt}}(t;X_{n+j})$ denote the IPCW p-value used to test the null hypothesis $H_{\mathrm{lt}}(t;j):T_{n+j}\geq t$, for $j=1,\dots,m$. Define the false discovery rate (FDR) as
@@ 4.2 Results
Table 2: Summary of screening performance for high- and low-risk patient selection methods based on different survival models, aggregated across different datasets, screening tasks, and repetitions. The screened proportion denotes the average fraction of test patients selected. Although false positives cannot be directly verified due to censoring, approximate verification is possible and classified as valid, dubious, or invalid.
@@ Acknowledgements
We thank the two anonymous referees for their detailed feedback and valuable suggestions on an earlier version of this manuscript.
@@ Appendix A Mathematical Proofs
Under the null hypothesis that $T_{n+j}\geq t$, because $\hat{s}_{\mathrm{lt}}(t;x)$ is monotone increasing in $t$ for all $x$, we have that, almost surely
$\hat{s}_{\mathrm{lt}}(t;X_{n+j})\leq\hat{s}_{\mathrm{lt}}(T_{n+j};X_{n+j}),$
@@
$\tilde{\phi}_{\mathrm{lt}}(t;X_{n+j})$ $:=\frac{1+\sum_{i=1}^{n}\mathbb{I}\left\{\hat{s}_{\mathrm{lt}}(T_{i};X_{i})\geq\hat{s}_{\mathrm{lt}}(t;X_{n+j})\right\}}{n+1}$
$\geq\frac{1+\sum_{i=1}^{n}\mathbb{I}\left\{\hat{s}_{\mathrm{lt}}(T_{i};X_{i})\geq\hat{s}_{\mathrm{lt}}(T_{n+j};X_{n+j})\right\}}{n+1}.$
@@
$\mathbb{P}\left[\tilde{\phi}_{\mathrm{lt}}(t;X_{n+j})\leq\alpha,\,T_{n+j}\leq t\right]$ $\leq\mathbb{P}\left[\frac{1+\sum_{i=1}^{n}\mathbb{I}\left\{\hat{s}_{\mathrm{lt}}(T_{i};X_{i})\geq\hat{s}_{\mathrm{lt}}(T_{n+j};X_{n+j})\right\}}{n+1}\leq\alpha\right]$
@@
where the last inequality follows immediately through standard conformal inference arguments from the fact that $(\hat{s}_{\mathrm{lt}}(T_{1};X_{1}),\ldots,\hat{s}_{\mathrm{lt}}(T_{n};X_{n}),\hat{s}_{\mathrm{lt}}(T_{n+j};X_{n+j}))$ are exchangeable.
@@
$p_{j}:=\hat{\phi}_{\mathrm{lt}}(t;X_{n+j}),\qquad p_{j}^{*}:=\phi^{*}_{\mathrm{lt}}(t;X_{n+j}):=\mathbb{P}\left(\hat{s}_{\mathrm{lt}}(T;X)\geq\hat{s}_{\mathrm{lt}}(t;X_{n+j})\right),$
@@
$p_{j}^{*}=\phi^{*}_{\mathrm{lt}}(t;X_{n+j})=1-F(\hat{s}_{\mathrm{lt}}(t;X_{n+j})),$
where $F$ is the CDF of the score variable $\hat{s}_{\mathrm{lt}}(T;X)$, with corresponding density $f=F^{\prime}$ uniformly bounded above and below across all $t>0$. Then, the change-of-variables formula gives that the density $f^{*}$ of $p^{*}_{j}$ is:
@@
which is bounded from both above and below by:
@@ A.3.1 Validity of Oracle p-Values
$\mathcal{E}_{t,\delta}:=\left\{\sup_{x\in\mathbb{R}^{d}}\left|\hat{\phi}_{\mathrm{lt}}(t;x)-\phi^{*}_{\mathrm{lt}}(t;x)\right|\leq\epsilon_{n}(\delta)\right\},$ $\epsilon_{n}(\delta):=\frac{1}{\omega_{\min}\pi}\left[\frac{2(\omega_{\min}+1)}{n}+8\sqrt{\frac{\log(2n/\delta)}{n}}\right]+2\Delta_{N}.$
@@
$\phi^{*}_{\mathrm{lt}}(t;x)$ $:=\mathbb{P}\left(\hat{s}_{\mathrm{lt}}(T;X)\geq\hat{s}_{\mathrm{lt}}(t;x)\right),$
$\phi^{*}_{\mathrm{rt}}(t;x)$ $:=\mathbb{P}\left(\hat{s}_{\mathrm{rt}}(T;X)\geq\hat{s}_{\mathrm{rt}}(t;x)\right),$
@@
$\hat{\phi}_{\mathrm{lt}}(t;x):=R_{n}\left(\hat{s}_{\mathrm{lt}}(t;x)\right),\quad\phi^{*}_{\mathrm{lt}}(t;x):=R\left(\hat{s}_{\mathrm{lt}}(t;x)\right).$
@@
$\sup_{x\in\mathbb{R}^{d}}\left|\hat{\phi}_{\mathrm{lt}}(t;x)-\phi^{*}_{\mathrm{lt}}(t;x)\right|$ $=\sup_{x\in\mathbb{R}^{d}}\left|R_{n}(\hat{s}_{\mathrm{lt}}(t;x))-R(\hat{s}_{\mathrm{lt}}(t;x))\right|$

### v3

@@ 1.2 Preview of Our Contributions
For example, in Figure 1, we illustrate a scenario where one wishes to identify “low-risk” patients—defined as those with more than an $p=80\%$ chance of surviving beyond time $t=3$ years. A patient is flagged if their conformal band lies entirely above the 80% line at $t=3$. Our method asymptotically guarantees that, among all flagged individuals, on average at least 80% will survive beyond this time. This guarantee holds pointwise over all fixed choices of the probability threshold $p$ and time horizon $t$, and for “low-risk” as well as “high-risk” screening, making the method flexible and broadly applicable.
@@ 2.2 Preliminaries: Review of Conformal Inference without Censoring
If complete (uncensored) data are available, this hypothesis can be tested as follows. Let $\hat{s}_{\mathrm{lt}}(t^{\prime};x,t)$ be any monotone increasing function of $t^{\prime}$, which may depend on $x\in\mathcal{X}$ and the fixed time $t\in\mathbb{R}_{+}$ in the definition of the null hypothesis (1). We refer to $\hat{s}_{\mathrm{lt}}$ as the left-tail nonconformity scoring function.
An intuitive choice of scoring function satisfying the required monotonicity property is
$\hat{s}_{\mathrm{lt}}(t^{\prime};x,t):=\left(1-\hat{S}_{T}(t^{\prime}\mid x)\right)\mathbb{I}\left[t^{\prime}\geq t\right],$ (2)
where $\hat{S}_{T}(t^{\prime}\mid x)\in(0,1)$ denotes the conditional survival probability at time $t^{\prime}$ for an individual with covariates $x$, estimated by the model ^$\hat{\mathcal{M}}_{T}$. We have found that this scoring function works well in practice, although it is not the only possible option.
Using any suitable monotone increasing scoring function $\hat{s}_{\mathrm{lt}}$, we compute a non-conformity score $\hat{s}_{\mathrm{lt}}(T_{i},X_{i},t)$ for each calibration point $i\in[n]$ as well as for the test point $X_{n+j}$, in the latter case using the placeholder value $t$ instead of the unknown event time $T_{n+j}$—i.e., defining the test score as $\hat{s}_{\mathrm{lt}}(t,X_{n+j},t)$. Then, following the approach of Jin and Candès (2023b), we define the left-tail conformal p-value as:
$\tilde{\phi}_{\mathrm{lt}}(t;X_{n+j}):=\frac{1+\sum_{i=1}^{n}\mathbb{I}\left\{\hat{s}_{\mathrm{lt}}(T_{i};X_{i},t)\geq\hat{s}_{\mathrm{lt}}(t;X_{n+j},t)\right\}}{n+1}.$ (3)
For example, in the special case where the scoring function takes the form defined in (2), under the relatively mild assumption that $\hat{S}_{T}(t\mid X_{n+j})<1$ almost surely, then it is easy to verify that the left-tail conformal p-value in (3) can be written as:
$\tilde{\phi}_{\mathrm{lt}}(t;X_{n+j}):=\frac{1+\sum_{i=1}^{n}\mathbb{I}\left[T_{i}\geq t\right]\mathbb{I}\left\{\hat{S}_{T}(T_{i}\mid X_{i})\leq\hat{S}_{T}(t\mid X_{n+j})\right\}}{n+1}.$
@@
Assume the calibration data and test point $(X_{n+j},T_{n+j})$ are exchangeable and, for a fixed $t>0$, the scoring function $\hat{s}_{\mathrm{lt}}(t^{\prime};x,t)$ is monotone increasing in $t^{\prime}$ for all $x$. Then, for any $\alpha\in(0,1)$, the conformal p-value $\tilde{\phi}_{\mathrm{lt}}(t;X_{n+j})$ defined in (3), computed using the full (uncensored) data, satisfies:
@@
which is useful for identifying low-risk individuals. In this case, we consider a right-tail scoring function assumed to be monotone decreasing in $t^{\prime}$; e.g.,
$\hat{s}_{\mathrm{rt}}(t^{\prime};x,t):=\hat{S}_{T}(t^{\prime}\mid x)\mathbb{I}\left[t^{\prime}\leq t\right],$ (5)
interpreted as the predicted probability of experiencing the event after time $t^{\prime}$ for an individual with covariates $x$. The corresponding right-tail conformal p-value is
$\tilde{\phi}_{\mathrm{rt}}(t;X_{n+j}):=\frac{1+\sum_{i=1}^{n}\mathbb{I}\left\{\hat{s}_{\mathrm{rt}}(T_{i};X_{i},t)\geq\hat{s}_{\mathrm{rt}}(t;X_{n+j},t)\right\}}{n+1},$
@@ 2.3 Conformal Inference for Survival Analysis under Right-Censoring
$\hat{\phi}_{\mathrm{lt}}(t;X_{n+j}):=\frac{1+\sum_{i=1}^{n}E_{i}\cdot\hat{w}(T_{i},X_{i})\cdot\mathbb{I}\left\{\hat{s}_{\mathrm{lt}}(T_{i};X_{i},t)\geq\hat{s}_{\mathrm{lt}}(t;X_{n+j},t)\right\}}{1+\sum_{i=1}^{n}E_{i}\cdot\hat{w}(T_{i},X_{i})},$ (7)
where the numerator includes only data points whose true event times are observed. For example, in the special case where the scoring function takes the form defined in (2), under the relatively mild assumption that $\hat{S}_{T}(t\mid X_{n+j})<1$ almost surely, the left-tail conformal p-value in (7) can be written as:
$\hat{\phi}_{\mathrm{lt}}(t;X_{n+j}):=\frac{1+\sum_{i=1}^{n}E_{i}\cdot\hat{w}(T_{i},X_{i})\cdot\mathbb{I}\left[T_{i}\geq t\right]\cdot\mathbb{I}\left\{\hat{S}_{T}(T_{i}\mid X_{i})\leq\hat{S}_{T}(t\mid X_{n+j})\right\}}{1+\sum_{i=1}^{n}E_{i}\cdot\hat{w}(T_{i},X_{i})}.$
@@
$\hat{\phi}_{\mathrm{rt}}(t;X_{n+j}):=\frac{1+\sum_{i=1}^{n}E_{i}\cdot\hat{w}(T_{i},X_{i})\cdot\mathbb{I}\left\{\hat{s}_{\mathrm{rt}}(T_{i};X_{i},t)\geq\hat{s}_{\mathrm{rt}}(t;X_{n+j},t)\right\}}{1+\sum_{i=1}^{n}E_{i}\cdot\hat{w}(T_{i},X_{i})}.$ (8)
In the special case where the scoring function takes the form defined in (5), under the relatively mild assumption that $\hat{S}_{T}(t\mid X_{n+j})>0$ almost surely, this can be written as:
$\hat{\phi}_{\mathrm{rt}}(t;X_{n+j}):=\frac{1+\sum_{i=1}^{n}E_{i}\cdot\hat{w}(T_{i},X_{i})\cdot\mathbb{I}\left[T_{i}\leq t\right]\cdot\mathbb{I}\left\{\hat{S}_{T}(T_{i}\mid X_{i})\geq\hat{S}_{T}(t\mid X_{n+j})\right\}}{1+\sum_{i=1}^{n}E_{i}\cdot\hat{w}(T_{i},X_{i})}.$
@@
    • a training set of cardinality $N$ used to estimate $\hat{w}(t;x)$, $\hat{s}_{\mathrm{lt}}(t^{\prime};x,t)$, and $\hat{s}_{\mathrm{rt}}(t^{\prime};x,t)$;
@@
  2. (A2)
    The scoring function $\hat{s}_{\mathrm{lt}}(t^{\prime};x,t)$ is almost surely monotone increasing in $t^{\prime}$ for all $x,t$, while $\hat{s}_{\mathrm{rt}}(t^{\prime};x,t)$ is monotone decreasing. Moreover, $\hat{s}_{\mathrm{lt}}(t;x,t)\in(0,1)$ and $\hat{s}_{\mathrm{rt}}(t;x,t)\in(0,1)$ for all $x,t$, almost surely.
@@
In Section 2.4, we will explain how to use IPCW conformal p-values to construct conformal survival bands.
@@ 2.4 Conformal Survival Bands for Screening High- or Low-Risk Individuals
4:  for each time $t\in\mathcal{T}$ do
5:     for each calibration point $(X_{i},T_{i})\in\mathcal{D}_{\text{cal}}^{\prime}$ do
6:        Compute $\hat{s}_{\mathrm{lt}}(T_{i};X_{i},t)$ and $\hat{s}_{\mathrm{rt}}(T_{i};X_{i},t)$ using Eqs. (2) and (5).
7:     end for
@@
9:        Compute scores $\hat{s}_{\mathrm{lt}}(t;X_{n+j},t)$ and $\hat{s}_{\mathrm{rt}}(t;X_{n+j},t)$ using Eqs. (2) and (5).
@@ 2.5 Theoretical Guarantees for Calibrated Survival Bands
    For all $t^{\prime}>0$, the score functions $\hat{s}_{\mathrm{lt}}(t^{\prime};X,t)$ admit a Lebesgue density $f_{t^{\prime}}$ satisfying
    $0<f_{\min}\leq f_{t^{\prime}}(u)\leq f_{\max}<\infty\quad\text{for all }u>0,$
@@
Under Assumptions (A1)(A5), fix a target level $\alpha\in(0,1)$ and survival cutoff time $t>0$. Let $p_{j}:=\hat{\phi}_{\mathrm{lt}}(t;X_{n+j})$ denote the IPCW p-value used to test the null hypothesis $H_{\mathrm{lt}}(t;j):T_{n+j}\geq t$, for $j=1,\dots,m$. Define the false discovery rate (FDR) as
@@ 4.2 Results
Table 2: Summary of screening performance for high- and low-risk patient selection methods based on different survival models, aggregated across datasets, screening tasks, and repetitions. The screened proportion denotes the average fraction of test patients selected. Although false positives cannot be directly verified due to censoring, approximate verification is possible and classified as valid, dubious, or invalid. Note that the aggregated screened proportion is always 0.5 for the Model and KM benchmarks, which do not account for uncertainty and always assign each patient to either the high- or low-risk group for each pair $(p,t)$ of screening parameters. In contrast, CSB may leave some patients unclassified—neither high- nor low-risk—when the $(p,t)$ point falls within the “gray area” uncertainty band (e.g., see Figure 1), resulting in lower and more variable aggregated screening rates.
@@ Acknowledgements
The authors thank two anonymous referees for detailed feedback and valuable suggestions on an earlier version of this manuscript. M. S. is partly supported by NSF grant DMS 2210637, a Capital One CREDIF Research Award, and a Google Research Scholar Award.
@@ Appendix A Mathematical Proofs
Under the null hypothesis that $T_{n+j}\geq t$, because $\hat{s}_{\mathrm{lt}}(t^{\prime};x,t)$ is monotone increasing in $t^{\prime}$ for all $x$, we have that, almost surely
$\hat{s}_{\mathrm{lt}}(t;X_{n+j},t)\leq\hat{s}_{\mathrm{lt}}(T_{n+j};X_{n+j},t),$
@@
$\tilde{\phi}_{\mathrm{lt}}(t;X_{n+j})$ $:=\frac{1+\sum_{i=1}^{n}\mathbb{I}\left\{\hat{s}_{\mathrm{lt}}(T_{i};X_{i},t)\geq\hat{s}_{\mathrm{lt}}(t;X_{n+j},t)\right\}}{n+1}$
$\geq\frac{1+\sum_{i=1}^{n}\mathbb{I}\left\{\hat{s}_{\mathrm{lt}}(T_{i};X_{i},t)\geq\hat{s}_{\mathrm{lt}}(T_{n+j};X_{n+j},t)\right\}}{n+1}.$
@@
$\mathbb{P}\left[\tilde{\phi}_{\mathrm{lt}}(t;X_{n+j})\leq\alpha,\,T_{n+j}\leq t\right]$ $\leq\mathbb{P}\left[\frac{1+\sum_{i=1}^{n}\mathbb{I}\left\{\hat{s}_{\mathrm{lt}}(T_{i};X_{i},t)\geq\hat{s}_{\mathrm{lt}}(T_{n+j};X_{n+j},t)\right\}}{n+1}\leq\alpha\right]$
@@
where the last inequality follows immediately through standard conformal inference arguments because $(\hat{s}_{\mathrm{lt}}(T_{1};X_{1},t),\ldots,\hat{s}_{\mathrm{lt}}(T_{n};X_{n},t),\hat{s}_{\mathrm{lt}}(T_{n+j};X_{n+j},t))$ are exchangeable.
@@
$p_{j}:=\hat{\phi}_{\mathrm{lt}}(t;X_{n+j}),\qquad p_{j}^{*}:=\phi^{*}_{\mathrm{lt}}(t;X_{n+j}):=\mathbb{P}\left(\hat{s}_{\mathrm{lt}}(T;X,t)\geq\hat{s}_{\mathrm{lt}}(t;X_{n+j},t)\right),$
@@
$p_{j}^{*}=\phi^{*}_{\mathrm{lt}}(t;X_{n+j})=1-F(\hat{s}_{\mathrm{lt}}(t;X_{n+j},t)),$
where $F$ is the CDF of the score variable $\hat{s}_{\mathrm{lt}}(T;X,t)$, with corresponding density $f(u)=F^{\prime}(u)$ uniformly bounded above and below across all $u>0$. Then, the change-of-variables formula gives that the density $f^{*}$ of $p^{*}_{j}$ is:
@@
which, thanks to Assumptions (A2) and (A5), we know to be bounded from both above and below by:
@@ A.3.1 Validity of Oracle p-Values
$\mathcal{E}_{t,\delta}$ :=$:=\left\{\sup_{x\in\mathbb{R}^{d}}\left|\hat{\phi}_{\mathrm{lt}}(t;x)-\phi^{*}_{\mathrm{lt}}(t;x)\right|\leq\epsilon_{n}(\delta)\right\},$
$\epsilon_{n}(\delta)$ $:=\frac{1}{\omega_{\min}\pi}\left[\frac{2(\omega_{\min}+1)}{n}+8\sqrt{\frac{\log(2n/\delta)}{n}}\right]+2\Delta_{N}.$
@@
$\phi^{*}_{\mathrm{lt}}(t;x)$ $:=\mathbb{P}\left(\hat{s}_{\mathrm{lt}}(T;X,t)\geq\hat{s}_{\mathrm{lt}}(t;x,t)\right),$
$\phi^{*}_{\mathrm{rt}}(t;x)$ $:=\mathbb{P}\left(\hat{s}_{\mathrm{rt}}(T;X,t)\geq\hat{s}_{\mathrm{rt}}(t;x,t)\right),$
@@
$\hat{\phi}_{\mathrm{lt}}(t;x):=R_{n}\left(\hat{s}_{\mathrm{lt}}(t;x,t)\right),\quad\phi^{*}_{\mathrm{lt}}(t;x):=R\left(\hat{s}_{\mathrm{lt}}(t;x,t)\right).$
@@
$\sup_{x\in\mathbb{R}^{d}}\left|\hat{\phi}_{\mathrm{lt}}(t;x)-\phi^{*}_{\mathrm{lt}}(t;x)\right|$ $=\sup_{x\in\mathbb{R}^{d}}\left|R_{n}(\hat{s}_{\mathrm{lt}}(t;x,t))-R(\hat{s}_{\mathrm{lt}}(t;x,t))\right|$